# Geometric spectral theory of quantum graphs

### v2

@@ 4 Advanced surgery
3rd proof of Theorem 2.1: “unfolding pendant edges” (Berkolaiko–K.–Kurasov–Mugnolo)
4.1 Definition.
@@
This is an example of transplantation, removing length (“mass”) from one part of the graph (which may be deleting or shortening edges) and attaching it elsewhere (inserting new edges or lengthening additional ones), usually in such a way as to preserve the total length. It is explored in detail in [12, Section 3.3].
4.2 Lemma.
@@
Keeping with our running intuition about the speed of diffusion, this kind of principle (and related principles involving “unfolding parallel edges”) says that “diffusion (in our usual $L^{2}$-sense) is slowed when connections in parallel are replaced by connections in sequence”. The elementary but somewhat delicate proof is given in [12, Proof of Theorem 3.18(4)].
Third proof of Theorem 2.1.
Fix a graph $\Gamma$. By Lemma 3.4 we may assume without loss of generality that $\Gamma$ is actually a tree, since cutting through any cycles can only lower $\mu_{2}$. Denote by $n=n(\Gamma)$ the number of leaves (degree one vertices) of $\Gamma$; $\Gamma$ is a path graph if and only if $n=2$. If $n\geq 3$, then $\Gamma$ must have at least one pair of neighbouring pendant edges. Create a new graph $\Gamma_{1}$ by unfolding these; then we have reduced the number of leaves by one, $n(\Gamma_{1})=n-1$, and $\mu_{2}(\Gamma_{1})\leq\mu_{2}(\Gamma)$ by Lemma 4.2. Repeating this argument inductively will produce a path graph, with lower $\mu_{2}$, after $n-2$ steps. ∎
@@
Interpolation between the theorems of Nicaise and Band–Lévy; 4th proof of Theorem 2.1 (Berkolaiko–K.–Kurasov–Mugnolo)
@@
for all compact graphs $\Gamma$ of total length $L$, equality being achieved (only) by path graphs $\simeq$ intervals; while the Theorem of Band–Lévy ([8, Theorem 2.1(2)], cf. Remark 3.2) states that
@@
for all doubly connected graphs $\Gamma$ of total length $L$, equality being achieved by loops $\simeq$ flat one-dimensional tori. [^3] There is actually a family of related minimisers owing to a basic surgery principle, but not relevant for our purposes here: we may glue the loop together at certain points without affecting the eigenvalue, if there is always an eigenfunction which takes on the same value at the points being glued. This is the refinement of Lemma 3.4 mentioned in the proof of Theorem 3.6. We can interpolate between the two inequalities based on the size of the doubly connected part of $\Gamma$:
4.3 Definition.
@@
4.4 Theorem.
[12, Theorem 6.3] Suppose $\Gamma$ has doubly connected part of total length $V\in[0,L]$. Let $\mathfrak{D}$ be the dumbbell graph (cf. Figure 6) of total length $L$, whose two loops have equal length $V/2$ each. Then
@@
Figure 7: Turning a graph (here an equilateral complete graph on four edges) into a pumpkin chain, based on the behaviour of an arbitrary but fixed second eigenfunction. $v_{-}$ and $v_{+}$ represent, respectively, where the eigenfunction reaches its minimum and maximum; its values at the other vertices are considered “critical”, and thus all points at which it attains those values (labelled in blue and red, respectively; electronic version in colour) are glued together. The resulting pumpkin chain has the same eigenvalue and eigenfunction (up to a canonical mapping between the graphs), such that the eigenfunction is now monotonically increasing along each path from $v_{-}$ to $v_{+}$. (Each of the sets of parallel edges betweek neighbouring vertices is a pumpkin.) Figure adapted from [12, Figure 5.1].
Step 2: The edges within each of the constituent pumpkins in the pumpkin chain from Step 1 may not have the same length. Apply a local symmetrisation argument (a local version of the symmetrisation method of Section 2) to make each of the pumpkins equilateral (although edges from different pumpkins are allowed to have different lengths). In Figure 7 this step is not necessary due to the symmetry of the original graph and its eigenfunction.
Step 3: Apply a transplantation argument to the symmetrised pumpkin chain to “push” the mass outwards to form a dumbbell. [^5] This step is actually quite complicated, and is performed individually on each of the two nodal domains of the pumpkin chain. While it is a transplantation, in practice it requires combining a transplantation argument with a cutting procedure and a Hadamard formula for the derivative of $\mu_{2}$ as a function of the edge lengths.

### v3

@@ 4.4 Advanced surgery
Third proof of Theorem 4.3: “unfolding pendant edges” (Berkolaiko–K.–Kurasov–Mugnolo)
Definition 4.29.
@@
This is an example of transplantation, removing length (“mass”) from one part of the graph (which may be deleting or shortening edges) and attaching it elsewhere (inserting new edges or lengthening additional ones), usually in such a way as to preserve the total length. It is explored in detail in [13, Section 3.3].
Lemma 4.30.
@@
Keeping with our running intuition about the speed of diffusion, this kind of principle (and related principles involving “unfolding parallel edges”) says that “diffusion (in our usual $L^{2}$-sense) is slowed when connections in parallel are replaced by connections in sequence”. The elementary but somewhat delicate proof is given in [13, Proof of Theorem 3.18(4)].
Proof 4.31 (Third proof of Theorem 4.3).
Fix a graph $\Gamma$. By Lemma 4.19 we may assume without loss of generality that $\Gamma$ is actually a tree, since cutting through any cycles can only lower $\mu_{2}$. Denote by $n=n(\Gamma)$ the number of leaves (degree one vertices) of $\Gamma$; $\Gamma$ is a path graph if and only if $n=2$. If $n\geq 3$, then $\Gamma$ must have at least one pair of neighbouring pendant edges. Create a new graph $\Gamma_{1}$ by unfolding these; then we have reduced the number of leaves by one, $n(\Gamma_{1})=n-1$, and $\mu_{2}(\Gamma_{1})\leq\mu_{2}(\Gamma)$ by Lemma 4.30. Repeating this argument inductively will produce a path graph, with lower $\mu_{2}$, after $n-2$ steps.
@@
Interpolation between the theorems of Nicaise and Band–Lévy; fourth proof of Theorem 4.3 (Berkolaiko–K.–Kurasov–Mugnolo)
@@
for all compact graphs $\Gamma$ of total length $L$, equality being achieved (only) by path graphs (equivalently, intervals); while the Theorem of Band–Lévy ([8, Theorem 2.1(2)], cf. Remark 4.17) states that
@@
for all doubly connected graphs $\Gamma$ of total length $L$, equality being achieved by loops, equivalently, flat one-dimensional tori. [^5] There is actually a family of related minimisers owing to a basic surgery principle, but not relevant for our purposes here: we may glue the loop together at certain points without affecting the eigenvalue, if there is always an eigenfunction which takes on the same value at the points being glued. This is the refinement of Lemma 4.19 mentioned in the proof of Theorem 4.23. We can interpolate between the two inequalities based on the size of the doubly connected part of $\Gamma$:
Definition 4.32.
@@
Theorem 4.33.
[13, Theorem 6.3] Suppose $\Gamma$ has doubly connected part of total length $V\in[0,L]$. Let $\mathfrak{D}$ be the dumbbell graph (cf. Figure 7) of total length $L$, whose two loops have equal length $V/2$ each. Then
@@
Step 2: The edges within each of the constituent pumpkins in the pumpkin chain from Step 1 may not have the same length. Apply a local symmetrisation argument (a local version of the symmetrisation method of Section 4.2) to make each of the pumpkins equilateral (although edges from different pumpkins are allowed to have different lengths). In Figure 8 this step is not necessary due to the symmetry of the original graph and its eigenfunction.
Step 3: Apply a transplantation argument to the symmetrised pumpkin chain to “push” the mass outwards to form a dumbbell. [^7] This step is actually quite complicated and is performed individually on each of the two nodal domains of the pumpkin chain. While it is a transplantation, in practice it requires combining a transplantation argument with a cutting procedure and a Hadamard formula for the derivative of $\mu_{2}$ as a function of the edge lengths.
@@
Figure 8: Turning a graph (here an equilateral complete graph on four edges) into a pumpkin chain, based on the behaviour of an arbitrary but fixed second eigenfunction. The vertices $v_{-}$ and $v_{+}$ represent, respectively, where the eigenfunction reaches its minimum and maximum; its values at the other vertices are considered “critical”, and thus all points at which it attains those values (labelled in blue and red, respectively) are glued together. The resulting pumpkin chain has the same eigenvalue and eigenfunction (up to a canonical mapping between the graphs), such that the eigenfunction is now monotonically increasing along each path from $v_{-}$ to $v_{+}$. (Each of the sets of parallel edges betweek neighbouring vertices is a pumpkin.) Figure adapted from [13, Figure 5.1].
Appendix: Exercises
  • 1.1
    Use the spectral theorem to prove Courant–Fischer (Theorem 2.4), and show that the minimum in (1a) is attained if and only if $u$ is an eigenvector corresponding to $\lambda_{1}$. Can you characterise equality in (1b) and (1c)?
  • 1.2
    Prove that if $u\in C^{1}(\overline{\Omega})$, then its classical partial derivative $\frac{\partial u}{\partial x_{j}}\in C(\overline{\Omega})$ is also a weak partial derivative of $u$. Conversely, prove that if $g_{j}$ is a weak partial derivative of $u$ and $g_{j}\in C(\overline{\Omega})$, then in fact $g_{j}$ is a classical partial derivative of $u$.
  • 1.3
    Prove Lemma 2.11.
  • 1.4
    Schrödinger operators. Given a bounded Lipschitz domain $\Omega\subset\mathbb{R}^{d}$, define a Hermitian form $a_{q}:H^{1}_{0}(\Omega)\times H^{1}_{0}(\Omega)\to\mathbb{C}$ for $q\in L^{\infty}(\Omega,\mathbb{R})$ by
    $a_{q}(u,v):=\int_{\Omega}\nabla u\cdot\overline{\nabla v}+qu\overline{v}\,\textrm{d}x.$
    • (a)
      Show that the form $a_{q}$ is bounded and $L^{2}$-elliptic on $H^{1}_{0}(\Omega)$, and find the operator $A_{q}$ on $L^{2}(\Omega)$ which is associated with $a_{q}$. Deduce that $A_{q}$ satisfies the conclusions of the spectral theorem, Theorem 2.3.
      Note: In particular, this proves Theorem 2.12(a), which corresponds to $q=0$.
    • (b)
      Use the variational characterisation of the eigenvalues to show that, given $q_{1},q_{2}\in L^{\infty}(\Omega,\mathbb{R})$, if $q_{1}\leq q_{2}$ almost everywhere in $\Omega$, then
      $\lambda_{k}(A_{q_{1}})\leq\lambda_{k}(A_{q_{2}})\qquad\text{for all }k\geq 1.$
  • 1.5
    The Robin Laplacian. Given a bounded Lipschitz domain $\Omega\subset\mathbb{R}^{d}$, for $\alpha\in\mathbb{R}$ define a form $a_{\alpha}:H^{1}(\Omega)\times H^{1}(\Omega)\to\mathbb{C}$ by
    $a_{\alpha}(u,v):=\int_{\Omega}\nabla u\cdot\overline{\nabla v}\,\textrm{d}x+\int_{\partial\Omega}\alpha u\overline{v}\,\textrm{d}s,$
    where the functions in the boundary integral are to be taken as the traces of $u$ and $\overline{v}$. Clearly $a_{\alpha}$ is Hermitian, and $a_{\alpha}$ coincides with the form $a$ from (8) when $\alpha=0$.
    • (a)
      Use the trace inequality (5) to prove that $a_{\alpha}$ is bounded and $L^{2}$-elliptic on $H^{1}(\Omega)$.
    • (b)
      Show that the operator $A_{\alpha}$ on $L^{2}(\Omega)$ associated with $a_{\alpha}$ is given by
      $D(A_{\alpha})$ =$=\left\{u\in H^{1}(\Omega):\Delta u\in L^{2}(\Omega),\,\frac{\partial u}{\partial\nu}\in L^{2}(\partial\Omega)\text{ and }=-\alpha u\text{ in }L^{2}(\partial\Omega)\right\},$
      $A_{\alpha}u$ $=-\Delta u,$
      and deduce that $A_{\alpha}$ satisfies the conclusions of the spectral theorem, Theorem 2.3.
  • 1.6
    The first eigenvalue of the Robin Laplacian $A_{\alpha}$ from the previous exercise may be characterised variationally as
    $\lambda_{1}(\alpha):=\lambda_{1}(A_{\alpha})=\min_{0\neq u\in H^{1}(\Omega)}\frac{\int_{\Omega}|\nabla u|^{2}\,\textrm{d}x+\int_{\partial\Omega}\alpha|u|^{2}\,\textrm{d}x}{\int_{\Omega}|u|^{2}\,\textrm{d}x}.$
    • (a)
      Prove that (on a fixed domain $\Omega$) the function $\alpha\mapsto\lambda_{1}(\alpha)$ is a convex monotonically increasing and continuous function of $\alpha\in\mathbb{R}$, and that $\lambda_{1}(\alpha)\leq\lambda_{1}(-\Delta^{D})$, the first eigenvalue of the Dirichlet Laplacian, for all $\alpha\in\mathbb{R}$.
    • (b)
      Now suppose $\alpha>0$.
      It follows immediately from (a) that $\lambda_{1}(\alpha)\searrow\lambda_{1}(-\Delta^{N})=0$, the first eigenvalue of the Neumann Laplacian, as $\alpha\to 0$. Show that $\lambda_{1}(\alpha)>0$ for all $\alpha>0$, and deduce that $\sqrt{a_{\alpha}(u,u)}$ defines an equivalent norm on $H^{1}(\Omega)$.
    • (c)
      Show that $\lambda_{1}(\alpha)\nearrow\lambda_{1}(-\Delta^{D})$ as $\alpha\to\infty$.
      Hint: Given $\alpha_{n}\to\infty$, let $\psi_{n}$ be an eigenfunction for $\lambda_{1}(\alpha_{n})$ with $L^{2}$-norm $1$. Deduce from (a) and (b) that $(\psi_{n})$ forms a bounded, and thus weakly convergent, sequence in $H^{1}(\Omega)$.
  • 2.1
    $W^{1,p}_{0}$-spaces. Given a (connected, compact) graph $\Gamma=(V,E)$, choose an arbitrary subset $V_{D}\subseteq V$ of the vertices of $\Gamma$ and define, for fixed $1\leq p\leq\infty$,
    $W^{1,p}_{0}(\Gamma):=W^{1,p}_{0}(\Gamma;V_{D}):=\{f\in W^{1,p}(\Gamma):f(v)=0\text{ for all }v\in V_{D}\}.$
    If $p=2$, then we usually write $H^{1}_{0}(\Gamma)=H^{1}_{0}(\Gamma;V_{D})$ in place of $W^{1,2}_{0}(\Gamma)=W^{1,2}_{0}(\Gamma;V_{D})$.
    • (a)
      Prove that $W^{1,p}_{0}(\Gamma;V_{D})$ is a closed subspace of $W^{1,p}(\Gamma)$.
      Hint: It follows from Lemma 3.13 that the embedding $W^{1,p}(\Gamma)\hookrightarrow C(\Gamma)$ is also continuous.
    • (b)
      Consider the $3$-star graph $\Gamma$ of Example 3.2. Show that if $V_{D}=\{v_{4}\}$, then $W^{1,p}_{0}(\Gamma;\{v_{4}\})$ is isometrically isomorphic to the direct sum of three copies of $W^{1,p}_{0}([0,1];\{1\})$. Thus, imposing a Dirichlet condition at a vertex of degree $\geq 2$ “decouples” the vertex.
    • (c)
      Still on the example of the $3$-star graph $\Gamma$, show that $W^{1,p}_{0}(\Gamma;\{v_{1},v_{2},v_{3}\})$ is not isometrically isomorphic to $W^{1,p}_{0}(\Gamma;\{v_{4}\})$. Obviously, the choice of the set $V_{D}$ matters!
    • (d)
      Returning to the general case, prove that the codimension of $H^{1}_{0}(\Gamma)$ in $H^{1}(\Gamma)$ is #$\#V_{D}$. Is the same true if $p\neq 2$?
  • 2.2
    Prove Lemma 3.20.
  • 2.3
    Prove Proposition 3.22 in the special case when $q=0$.
  • 2.4
    The Laplacian with (some) Dirichlet vertex conditions. Given a (connected, compact) graph $\Gamma=(V,E)$, choose a set $V_{D}\subseteq V$ and consider the form
    $a_{0}(f,g)=\int_{\Gamma}f^{\prime}\overline{g^{\prime}}\,\textrm{d}x$
    from (13), but this time defined on $H^{1}_{0}(\Gamma;V_{D})$ instead of $H^{1}(\Gamma)$. It is obvious that Lemma 3.20 continues to hold in this space, and thus the spectral theorem holds for the associated operator.
    • (a)
      Prove that the associated operator $A_{0}^{D}$ is given by
      $D(A_{0}^{D})$ =$=\big{\{}f\in H^{1}(\Gamma):f^{\prime\prime}\in L^{2}(\Gamma),\,\,f(v)=0\,\,\forall v\in V_{D},$
[…]
      $A_{0}^{D}f$ $=-f^{\prime\prime}$
      in the sense of distributions. Thus, functions in $D(A_{0}^{D})$ satisfy a Dirichlet condition at every vertex in $V_{D}$ and standard conditions at all others.
    • (b)
      Prove that, for any $V_{D}\subseteq V$,
      $\lambda_{k}(A_{0})\leq\lambda_{k}(A_{0}^{D})\leq\lambda_{k+(\#V_{D})}(A_{0}),$
      where the $\lambda_{k}(A_{0})$ are the eigenvalues of the standard Laplacian.
      Hint: Use the results of 2.1(a) and (d) together with the variational (min-max) characterisation of the eigenvalues.
    • (c)
      Prove that if $V_{D}\neq\emptyset$, then $\lambda_{1}(A_{0}^{D})>0$.
    • (d)
      Repeat this exercise (with appropriate modifications) for the following operator. Given a set of vertices $V_{R}\subseteq V$ and $\alpha>0$, the form is taken to be
      $a_{\alpha}(f,g)=\int_{\Gamma}f^{\prime}\overline{g^{\prime}}\,\textrm{d}x+\sum_{v\in V_{R}}\alpha f(v)\overline{g(v)}$
      on $H^{1}(\Gamma)$.
      Note: The associated operator is the Laplacian with Robin, a.k.a. $\delta$, vertex conditions on $V_{R}$ (and standard conditions elsewhere). The constant $\alpha$, which may also be allowed to depend on the vertex, represents a $\delta$-potential being imposed at the vertices $V_{R}$. This constant may also be taken non-positive, but then the conclusions of (b) and (c) are not valid. The eigenvalues do, however, satisfy the same conclusions as in 1.6.
  • 2.5
    The discrete Laplacian on the $3$-star graph. We will study a prototypical discrete graph which is the analogue of the $3$-star graph of Example 3.2.
    • (a)
      With the vertex numbering of Example 3.2 (in particular, where $\deg v_{4}=3$), write down the adjacency matrix $A$ of $\Gamma$, that is, $A_{ij}=1$ if there is an edge between $v_{i}$ and $v_{j}$, and 00 otherwise.
    • (b)
      Let $D$ be the $4\times 4$-diagonal matrix whose $(i,i)$-th entry is the degree of the vertex $v_{i}$, and let
      $L=D-A.$
      Determine $L$, and determine its spectrum. $L$ is the discrete Laplacian on $\Gamma$. (Observe that $L$ is a symmetric, real-valued matrix.)
    • (c)
      Show that a vector $u\in\mathbb{C}^{4}$ satisfies $Lu=0$ if and only if $u$ has the mean value property: for every vertex $v_{k}$,
      $u(v_{k})=\frac{1}{\deg v_{k}}\sum_{v_{j}\sim v_{k}}u(v_{j}).$
    These definitions and properties hold for the discrete Laplacian on any finite graph $\Gamma$.
  • 3.1
    Let $\Gamma_{m}$ be the (equilateral) $m$-star graph consisting of $m$ edges of length $\frac{1}{m}$ each, attached together at a vertex of degree $m$. (That is, formally, $E=\{e_{1},\ldots,e_{m}\}$, $V=\{v_{1},\ldots,v_{m+1}\}$, with $e_{k}\sim v_{k}v_{m+1}$, where $\deg v_{k}=1$ for $k=1,\ldots,m$ and $\deg v_{m+1}=m$.)
    • (a)
      Determine the lowest nontrivial eigenvalue $\mu_{2}(\Gamma_{m})$ of the standard Laplacian on $\Gamma_{m}$, and conclude that $\mu_{2}(\Gamma_{m})\to\infty$ even though $|\Gamma_{m}|=1$ for all $m\geq 3$.
    • (b)
      Describe all eigenvalues and eigenfunctions of $\Gamma_{m}$. Take care with the high multiplicities.
  • 3.2
    Use the characterisation of equality in (16) to prove the following special case of equality in Lemma 4.19: Suppose $\Gamma$ is formed from $\widetilde{\Gamma}$ by gluing together two vertices $v_{1}$ and $v_{2}$, and suppose also that there is an eigenfunction $\psi_{2}$ for $\mu_{2}(\widetilde{\Gamma})$ such that $\psi_{2}(v_{1})=\psi_{2}(v_{2})$. Then $\mu_{2}(\Gamma)=\mu_{2}(\widetilde{\Gamma})$.
  • 3.3
    Weyl asymptotics. In this exercise we will prove Corollary 4.21, that if $\Gamma$ has total length $L$, then $\mu_{k}(\Gamma)=\frac{\pi^{2}k^{2}}{L^{2}}+o(k^{2})$ as $k\to\infty$.
    • (a)
      Prove that the assertion holds if $\Gamma$ is just an interval of length $L$.
    • (b)
      Use (a) to show that the assertion also holds if $\Gamma$ is a disjoint union of $m$ intervals (of lengths $\ell_{1},\ldots,\ell_{m}>0$).
      Hint: It may be more practical to work with the counting function
      $N(\lambda):=\#\{k\in\mathbb{N}:\mu_{k}(\Gamma)\leq\lambda\}.$
    • (c)
      Use (b) and (22) to obtain the Weyl asymptotics for an arbitrary compact metric graph $\Gamma$.
    • (d)
      Prove that the Weyl asymptotics is still valid if any mix of standard, Dirichlet and Robin conditions (with $\alpha>0$ for simplicity) is imposed at the vertices.
  • 3.4
    • (a)
      Prove Rohleder’s inequality [44, Theorem 3.4]: Suppose $\Gamma$ is a tree graph, i.e. a graph with no cycles, and define its diameter to be the longest path in the graph,
      $D:=\max\{\operatorname{dist}(x,y):x,y\in\Gamma\}.$
      Then, for any $k\geq 1$,
      $\mu_{k}(\Gamma)\leq\frac{\pi^{2}k^{2}}{D^{2}}=\mu_{k}(\text{interval of length$D$}).$
    • (b)
      (Hard.) Does the same inequality hold if $\Gamma$ is not a tree graph?
    • (c)
      (Less hard but still hard.) Prove that no corresponding lower bound is possible. That is, find a sequence of tree graphs $\Gamma_{n}$ such that each has diameter $1$, but $\mu_{2}(\Gamma_{n})\to 0$ as $n\to\infty$.